When I solve the following problem to complete the square I find no solution to the problem is this correct or (good chance) I did something wrong. I am so confused.
3x^2-2x+5=0
Please help.
Thanks

Answer by venugopalramana(3286)

YOU ARE CORRECT. HAVE CONFIDENCE..BELIEVE IN YOUR SELF..THE PROBLEM HAS NO REAL NUMBER SOLUTION
ITS DISCRIMINANT =(-2)^2-4*3*5=4-60=-56... IS NEGATIVE.SO IT HAS NO REAL SOLUTION.
ON COMPLETING THE SQUARE YOU GET NEGATIVE VALUE FOR A PERFECT SQUARE WHICH IS NOT POSSIBLE IN REAL NUMBERS